OpenStudy (anonymous):

This totally confused me!! Find the first 5 terms of the sequence: a1 = 500, an = (an-1)/5.

OpenStudy (across):

This is a recursive definition of a sequence.\[a_1=500,\]\[a_n=\frac{a_{n-1}}{5}.\]The first term is trivially \(500\). The second one is\[a_2=\frac{a_1}{5}=\frac{500}{5}=100,\]the third one is\[a_3=\frac{a_2}{5}=\frac{100}{5}=20,\]and so on.

OpenStudy (anonymous):

ok(: Thank you so much. You've been a huge help